# On Singular Integrals and Quantitative Rectifiability in Parabolic Space and the Heisenberg Group

### v1

@@ Abstract.
David and Semmes proved that if all CZOs (of suitable dimension) are bounded with respect to an Ahlfors regular measure, then the measure is uniformly rectifiable. We extend this theorem to the parabolic space and the first Heisenberg group.
@@ 1. Introduction
Understanding the geometry of a measure in $\mathbb{R}^{n}$ for which an associated singular integral operator is well behaved is a central problem in harmonic analysis, with its origins in understanding removable sets for analytic capacity [Gar, Tol2, Tol3, Dav1, DM, NTV1, NTV2, NTV3] and boundary value problems for elliptic PDE [AMT, AHMTV, HM, HMM, HMU]. David and Semmes introduced the notion of uniform rectifiability, and discovered that it was a natural condition from the standpoint of singular integral operators:
@@
Suppose that $\mu$ is a $k$-Ahlfors regular (henceforth $k$-ADR) for $k\in\{1,\dots,n-1\}$. Then every convolution singular integral operator associated with a kernel $K\in C^{\infty}(\mathbb{R}^{n}\backslash\{0\})$ satisfying
@@
Here $\beta_{\mu}(Y,r)$ measures the deviation of the measure $\mu$ from an appropriate $k$-dimensional plane which optimally approximates (in the least squares sense) $\operatorname{supp}(\mu)$ in the ball $B(Y,r)$. In [DS1] they prove the ‘only if’ direction of Theorem 1.1 (from the boundedness of CZOs to the uniform rectifiability), and it is this direction that we will refer to as the David-Semmes theorem in this paper. On the other hand, the ‘if’ direction of Theorem 1.1 is a culmination of foundational work studying Calderón-Zygmund operators on Lipschitz graphs by Calderón and Coifman–McIntosh–Meyer.
@@
David and Semmes asked whether $L^{2}(\mu)$-boundedness of the $k$-Riesz transform [^1] The CZO with kernel $K(x)=\frac{x}{|x|^{k+1}}$ for $x\in\mathbb{R}^{n}$. alone implies uniform rectifiability. This question has been answered affirmatively in the case $k=1$ by Mattila-Melnikov-Verdera [MMV], and k=n1k=n-1 by Nazarov-Tolsa-Volberg [NToV]. These results, and refinements of them, have played an essential role in recent breakthroughs in the study of harmonic measure and free boundary problems (described later). The David-Semmes question remains open for $k=2,\dots,n-2$. The Nazarov-Tolsa-Volberg theorem [NToV] has been extended to “Riesz transforms” associated with gradients of fundamental solutions to more general classes of elliptic equations. Prat, Puliatti, and Tolsa extended the Nazarov-Tolsa-Volberg result to elliptic equations associated to matrices with Hölder coefficients. The current state-of-the-art is due to Molero, Mourgoglou, Puliatti, and Tolsa in [MMPT], where they prove the Nazarov-Tolsa-Volberg result for elliptic equations associated to matrices with coefficients satisfying a Dini-mean oscillation condition. Dabrowski and Tolsa have recently extended the Nazarov-Tolsa-Volberg theorem to measures with significantly weaker density hypotheses than ADR in a very significant work [DT].
@@
Some very important machinery in the proof of the Nazarov-Tolsa-Volberg result was introduced by Eiderman, Nazarov, and Volberg in [ENV], where the authors show that the Riesz transform cannot be $L^{2}$-bounded with respect to a measure with poor lower-density properties. The result of [ENV] was later extended to “Riesz transforms” associated to elliptic equations with Hölder coefficients by Conde-Alonso, Mourgoglou, and Tolsa in [CMT].
@@
In a recent survey on rectifiability, Mattila highlights the scarcity of results concerning the geometric consequences of the boundedness of singular integrals in parabolic space and in the Heisenberg group (see Sections 11.5 and 11.6 of [Mat]). The goal of this paper is to take a step in this direction by developing David–Semmes-type theorems in these settings. In the Heisenberg group, we show that a result analogous to Theorem 1.1 holds. In contrast, we find that parabolic space behaves quite differently: the analogue of Theorem 1.1 in parabolic space is, in fact, false. Although we provide results in all integer codimensions, we first describe the parabolic codimension-one case here; the higher-codimension case is more intricate and will be discussed later (Section 1.3). Specifically, there exist codimension-one Ahlfors regular measures in parabolic space for which a large class of parabolic Calderón–Zygmund operators (CZOs) are $L^{2}$-bounded with respect to the measure, yet the measures fail to be parabolically uniformly rectifiable. The “large class of CZOs” is defined precisely in Subsection 1.1. In particular, this class is designed to retain the essential features of the parabolic Riesz transform—that is, the parabolic CZO obtained by convolving with the spatial gradient of the fundamental solution to the heat equation. These counterexample measures possess a peculiar structure: they are highly concentrated in the spatial component of parabolic space. We recover a David–Semmes-type theorem in the parabolic setting by imposing, in addition to the boundedness of singular integrals, a new Carleson set condition called the HSDC (higher spatial dimension condition). This condition asserts that it is rare for a measure to be close, in the sense of optimal transport distance, to one of these pathological examples. For codimension-one measures in parabolic space, we prove the following theorem (see Section 1.3 for the precise statement).
@@
Mattila ([Mat]) has identified that understanding the consequences of boundedness of the parabolic Riesz transform as an important step in understanding the relationship between absolute continuity of caloric measure and parabolic rectifiability. The examples we construct show that the naïve translation of the Nazarov-Tolsa-Volberg theorem to the parabolic setting is false, while Theorem 1.2 gives us hope that a sharp theorem can still be recovered through the HSDC condition. We conjecture the following.
@@
In the case of a parabolic, codimension-one Ahlfors regular measure $\mu$, the universally agreed upon definition of uniform rectifiability is estimate (1.1), with $\beta_{\mu}$ measuring the “least squares” distance to an optimal vertical hyperplane, i.e. a codimension-one hyperplane containing a line parallel to the $t$-axis (see Section 1.2). This definition of parabolic uniform rectifiability was introduced by Hofmann, Lewis, and Nyström in [HLN]. To our knowledge, there has been no prior work about parabolic uniform rectifiability in codimension greater than one. In the sequel we will define parabolic uniformly rectifiable measures on $\mathbb{R}^{n+1}$ of higher codimension as Ahlfors regular measures (with appropriate homogeneity) satisfying a version of (1.1) where the $\beta_{\mu}$ coefficients optimize the distance to a higher-codimensional vertical plane.
@@
Suppose $\mu$ is a $3$-regular measure on $\mathbb{H}$ on which Calderón-Zygmund singular integral operators of dimension $3$ are bounded from $L^{2}(\mu)$ to $L^{2}(\mu)$. Then $\mu$ is a uniformly rectifiable measure (i.e. (1.2) holds).
@@
As far as we are aware, there is no standard definition of “uniform rectifiability” in the Heisenberg group literature. In analogy with the definition in parabolic space, we say that a $3$-ADR measure is uniformly rectifiable if it satisfies a version of estimate (1.1) where $\beta_{\mu}$ measures the distance to an optimal vertical hyperplane (Section 1.2). Recent works by Chousionis, Li, and Young provide context for Theorem 1.4 and the importance of the condition (1.1) in the first Heisenberg group. In [CLY1], the authors construct an intrinsic Lipschitz graph (i.e., a graph which is defined with respect to the exterior cone condition in the Heisenberg metric) on which the Riesz transforms associated to fundamental solution of the sub-Laplacian is not $L^{2}(\mu)$ bounded. They also show that the Lipschitz constant of a given intrinsic Lipschitz graph cannot control the constant of the right-hand side of inequality (1.1) with the exponent $2$ replaced by any $p\in[2,4)$. In [CLY3], they show that any given intrinsic Lipschitz graph in $\mathbb{H}$ satisfies (1.1) with exponent $2$ replaced by $4$, with the constant on the right-hand-side depending only the Lipschitz constant of the graph. Both of these results are proved using techniques developed in [NY2], which we discuss below. This is in contrast to higher Heisenberg groups $\mathbb{H}^{n}$, where, in [CLY2], the same authors show that intrinsic Lipschitz graphs automatically satisfy an estimate like (1.1). Thus, in the presence of the Ahlfors regularity condition, it is natural to seek “$1$-Heisenberg variants” of (1.1) from hypotheses involving boundedness of singular integrals.
The behavior of singular integrals on the Heisenberg group have been heavily studied, with important foundational work by Chousionis and Mattila [CM]. Significant results regarding sufficient conditions for the boundedness of all CZOs have been developed in Chousionis, Fässler and Orponen [CFO1, FO]. In [CLY1, CLY2], Chousionis, Li and Young conjecture that uniform rectifiability is sufficient for the boundedness of all CZOs (Theorem 1.4 provides the necessity) – we do not consider this interesting question here.
In terms of prior work concerning Theorem 1.4, Orponen [Orp] had previously verified a weaker conclusion that for every $\mathbf{X}\in\operatorname{supp}(\mu)$ and $R>0$,
$\int_{0}^{R}\mathcal{H}^{3}(\{\mathbf{Y}\in\operatorname{supp}(\mu):\beta_{\mu}(B(\mathbf{Y},r))\geq\varepsilon\})\frac{dr}{r}\leq CR^{3}.$
@@
The Nazarov-Tolsa-Volberg theorem ([NToV]) has had profound consequences for the study of free boundary problems involving elliptic measures, and, relatedly, connections between continuity of elliptic measures and rectifiability (both qualitative and quantitative). We hope that our results will be an important first step toward establishing analogs of the Nazarov-Tolsa-Volberg theorem in parabolic space and the Heisenberg group. We also think that the HSDC condition should play an important role in the theory of caloric and parabolic measures and parabolic rectifiability.
@@
In 1992 Bishop ([Bi, Conjecture 10]) conjectured that absolute continuity of harmonic measure with respect to surface measure should imply rectifiability of the boundary. Bishop’s conjecture was solved in 2015 in [AHMTV] where the authors effectively used estimates on the Green function to control the Riesz transform and verify the hypotheses of the codimension-one David-Semmes conjecture. Later, in [MT], Mourgoglou and Tolsa quantified the main result of [AHMTV] using extremely technical arguments involving a corona decomposition with respect to harmonic measure.
@@
“Two-phase” free boundary problems for harmonic (or, more generally, elliptic measures) are problems where one assumes that two harmonic measures in distinct connected components of a domain enjoy some mutual absolute continuity, in order to conclude that the boundary of the domain enjoys some regularity. In [Bi], Bishop also conjectured that mutual absolute continuity of harmonic measures in separated connected components of a domain should imply rectifiability of the common boundary of those components. Two-phase problems are more delicate than one phase problems, but similar (and more difficult) arguments can be used to tackle them. In particular, Girela-Sarrión and Tolsa proved a more delicate version of the codimension-one David-Semmes problem in [GT] which is suitable for applications to two-phase problems. The main result of [GT] was later used in solutions to qualitative and quantitative two-phase problems harmonic measure, we refer the reader to [AMT], [AMT2], and [AMTV] (the first two authored by Azzam, Mourgoglou, and Tolsa, and the last authored by Azzam, Mourgoglou, Tolsa, and Volberg). In particular, [AMTV] resolves the “two-phase” conjecture of Bishop ([Bi]). Very recently, Merlo, Mourgoglou, and Puliatti ([MMP]) have extended many of the results above (one-phase problems, two-phase problems, and the result of Girela-Sarrión and Tolsa) to elliptic measures with coefficients satisfying a “Dini-mean oscillation” condition. We also record that, prior to [MMP], Puliatti ([P]) had extended the result of Girela-Sarrión and Tolsa to elliptic equations associated to matrices with Hölder coefficients.
@@
In the parabolic case, a converse to Theorem 1.2 is known. In [BHHLN1, BHHLN2, BHHLN3] Bortz, Hofmann, Luna-García and Nyström, together with the first author, developed several foundational quantitative tools for the theory of parabolic quantitative rectifiability. In particular, in [BHHLN2] and [BHHLN3] together imply the following
@@
Suppose $\mu$ is a codimension-one parabolic uniformly rectifiable measure on $\mathbb{R}^{n+1}$. Then, if $T$ is a singular integral operator associated to a kernel $K:\mathbb{R}^{n+1}\setminus\{0\}\rightarrow\mathbb{R}$ such that
@@
We remark that graphs of parabolic Lipschitz functions need not be parabolic uniformly rectifiable (see [BHHLN1] for a counter-example). This is in contrast to the Euclidean setting, where Lipschitz graphs are a prototypical example of a uniformly rectifiable set. Very recently, the authors have extended their result to the case when $\Gamma$ is a time-symmetric ADR set [^2] Time-symmetric ADR is a stronger condition than the $\bf HSDC$}, see Corollary 1.9. The time symmetric ADR condition is a condition that forces every surface ball to have ample intersection with the set in forward and backward in time.
Theorem 1.7 ([BHMN2]).
@@
Neither proof relies on analysis of the Riesz transform. In analogy with the elliptic case, we expect that a positive resolution of Conjecture 1.3 would lead to further extensions and refinements of these results.
In the past decade there have been some extraordinary results concerning quantitative rectifiability and Heisenberg groups. In [NY1], Naor and Young resolved several open problems by proving a discrete “vertical versus horizontal” isoperimetric inequality in a discrete Heisenberg group $\mathbb{H}^{n}_{\mathbb{Z}}$, $n\geq 2$. In order to prove this inequality, they prove a corresponding inequality in the continuous Heisenberg group $\mathbb{H}^{n}$, and use the continuous case to deduce the discrete case. The methods of their proof utilize many tools from the area of quantitative rectifiability, in particular the “corona decomposition” type arguments introduced by David and Semmes in [DS1]. This discrete isoperimetric inequality is used to find an optimal (up to a constant) lower bound on the integrality gap of the Goemans–Linial semidefinite program for the Sparsest Cut Problem, and also establish a lower bound on the distortion of any embedding from the ball of radius $r$ centered at the origin in $(\mathbb{H}_{\mathbb{Z}}^{2},d_{W})$ (the second discrete Heisenberg group equipped with the “word metric”) is at least $C\sqrt{r}$. In [NY1], Naor and Young prove a “vertical-versus-horizontal” Poincaré inequality in the first Heisenberg group. As a corollary, they obtain, in the first Heisenberg group, bounds on the distortion of embeddings of balls centered at the origin. In this work they develop a new multiscale decomposition for intrinsic Lipschitz graphs, which was later modified in the aforementioned works [CLY1, CLY2].
There has also been significant advances in extending Preiss’ theorem to the Heisenberg group and parabolic space. These have been developed in the codimension-one setting by Merlo ([M]), and Merlo, Mourgoglou, and Puliatti ([MMP]), respectively.

### v2

@@ Abstract.
David and Semmes proved that if a large class of Calderón–Zygmund singular integral operators (of suitable homogeneity) are bounded with respect to an Ahlfors regular measure, then the measure is uniformly rectifiable. We extend this theorem to parabolic space and the first Heisenberg group. Notably, we show that the direct analog in parabolic space is false by way of counterexamples. However, we recover a theorem in parabolic space by imposing a “Carleson packing” condition, which stipulates that most scales of the measure are far away from the aforementioned counterexamples, in the sense of Tolsa’s $\alpha$-numbers.” In the Heisenberg group, we leverage previous work by Orponen to show that no such counterexamples exist, and hence we obtain a direct analog of the David–Semmes theorem in the first Heisenberg group. Our counterexamples in parabolic space demonstrate that there exist ADR measures on which the parabolic Riesz transform (associated to the gradient of the fundamental solution of the heat equation) is $L^{2}$-bounded, but which are not parabolic uniformly rectifiable.
1991 Mathematics Subject Classification:
28,31,35,42,44,51
@@ 1. Introduction
Understanding the geometry of a measure in $\mathbb{R}^{n}$ for which an associated singular integral operator is well behaved is a central problem in harmonic analysis, with its origins in understanding removable sets for analytic capacity [27, 54, 55, 22, 19, 49, 48, 50] and boundary value problems for elliptic PDE [3, 1, 32, 30, 31]. David and Semmes introduced the notion of uniform rectifiability, and discovered that it was a natural condition from the standpoint of singular integral operators:
@@
Suppose that $\mu$ is a $k$-Ahlfors regular (henceforth $k$-ADR) on $\mathbb{R}^{n}$ for $k\in\{1,\dots,n-1\}$. Then every convolution singular integral operator associated with a kernel $K\in C^{\infty}(\mathbb{R}^{n}\backslash\{0\})$ satisfying
@@
Here $\beta_{\mu}(\mathbf{Y},r)$ measures the deviation of the measure $\mu$ from an appropriate $k$-dimensional plane which optimally approximates (in the least squares sense) $\operatorname{supp}(\mu)$ in the ball $B(Y,r)$. In [20] they prove the ‘only if’ direction of Theorem 1.1 (from the boundedness of CZOs to the uniform rectifiability), and it is this direction that we will refer to as the David-Semmes theorem in this paper. On the other hand, the ‘if’ direction of Theorem 1.1 is a culmination of foundational work studying Calderón-Zygmund operators on Lipschitz graphs by Calderón and Coifman–McIntosh–Meyer.
@@
David and Semmes asked whether $L^{2}(\mu)$-boundedness of the $k$-Riesz transform [^1] The CZO with kernel $K(x)=\frac{x}{|x|^{k+1}}$ for $x\in\mathbb{R}^{n}$. alone implies uniform rectifiability. This question has been answered affirmatively in the case $k=1$ by Mattila-Melnikov-Verdera [39], and k=n1k=n-1 by Nazarov-Tolsa-Volberg [47]. These results, and refinements of them, have played an essential role in recent breakthroughs in the study of harmonic measure and free boundary problems (described later). The David-Semmes question remains open for $k=2,\dots,n-2$. The Nazarov-Tolsa-Volberg theorem [47] has been extended to “Riesz transforms” associated with gradients of fundamental solutions to more general classes of elliptic equations. For example, Prat, Puliatti, and Tolsa extended the Nazarov-Tolsa-Volberg result to elliptic equations associated to matrices with Hölder coefficients in [52]. The current state-of-the-art is due to Molero, Mourgoglou, Puliatti, and Tolsa in [43], where they prove the Nazarov-Tolsa-Volberg result for elliptic equations associated to matrices with coefficients satisfying a Dini-mean oscillation condition. Dabrowski and Tolsa have recently extended the Nazarov-Tolsa-Volberg theorem to measures with significantly weaker density hypotheses than ADR in a very significant work [18].
@@
Some very important machinery in the proof of the Nazarov-Tolsa-Volberg result was introduced by Eiderman, Nazarov, and Volberg in [24], where the authors show that the Riesz transform cannot be $L^{2}$-bounded with respect to a measure with poor lower-density properties. The result of [24] was later extended to “Riesz transforms” associated to elliptic equations with Hölder coefficients by Conde-Alonso, Mourgoglou, and Tolsa in [17]. In [43], Molero, Mourgoglou, Puliatti, and Tolsa extended the result further to operators with DMO-type (Dini mean oscillation) coefficients.
@@
In a recent survey on rectifiability, Mattila highlights the scarcity of results concerning the geometric consequences of the boundedness of singular integrals in parabolic space and in the Heisenberg group (see Sections 11.5 and 11.6 of [40]). The goal of this paper is to take a step in this direction by developing David–Semmes-type theorems in these settings. In the Heisenberg group, we show that a result analogous to Theorem 1.1 holds. In contrast, we find that parabolic space behaves quite differently: the analogue of Theorem 1.1 in parabolic space is, in fact, false. Although we provide results in all integer codimensions, we first describe the parabolic codimension-one case here; the higher-codimension case is more intricate and will be discussed later (Section 1.3). Specifically, there exist codimension-one Ahlfors regular measures in parabolic space for which a large class of parabolic Calderón–Zygmund operators (CZOs) are $L^{2}$-bounded with respect to the measure, yet the measures fail to be parabolically uniformly rectifiable. The “large class of CZOs” is defined precisely in Subsection 1.1. In particular, this class is designed to retain the essential features of the parabolic Riesz transform—that is, the parabolic CZO obtained by convolving with the spatial gradient of the fundamental solution to the heat equation. These counterexample measures possess a peculiar structure: they are highly concentrated in the spatial component of parabolic space. We recover a David–Semmes-type theorem in the parabolic setting by imposing, in addition to the boundedness of singular integrals, a new Carleson set condition called the HSDC (higher spatial dimension condition). This condition asserts that it is rare for a measure to be close, in the sense of optimal transport distance, to one of these pathological examples. For codimension-one measures in parabolic space, we prove the following theorem (see Section 1.3 for the precise statement).
@@
Mattila [40] has emphasized that understanding the consequences of the boundedness of the parabolic Riesz transform is a key step toward clarifying the relationship between absolute continuity of caloric measure and parabolic rectifiability. The examples we construct show that the naïve translation of the Nazarov-Tolsa-Volberg theorem to the parabolic setting is false, while Theorem 1.2 gives us hope that a sharp theorem can still be recovered through the HSDC condition. We conjecture the following.
@@
In the case of a parabolic, codimension-one Ahlfors regular measure $\mu$, the universally agreed upon definition of uniform rectifiability is estimate (1.1), with $\beta_{\mu}$ measuring the “least squares” distance to an optimal vertical hyperplane, i.e. a codimension-one hyperplane containing a line parallel to the $t$-axis (see Section 1.2). This definition of parabolic uniform rectifiability was introduced by Hofmann, Lewis, and Nyström in [29]. To our knowledge, there has been no prior work about parabolic uniform rectifiability in codimension greater than one. In the sequel we will define parabolic uniformly rectifiable measures on $\mathbb{R}^{n+1}$ of higher codimension as Ahlfors regular measures (with appropriate homogeneity) satisfying a version of (1.1) where the $\beta_{\mu}$ coefficients optimize the distance to a higher-codimensional vertical plane.
@@
Suppose $\mu$ is a $3$-ADR measure on $\mathbb{H}$ on which Calderón-Zygmund singular integral operators of dimension $3$ are bounded from $L^{2}(\mu)$ to $L^{2}(\mu)$. Then $\mu$ is a uniformly rectifiable measure (i.e. (1.2) holds).
@@
As far as we are aware, there is no standard definition of “uniform rectifiability” in the Heisenberg group literature. In analogy with the definition in parabolic space, we say that a $3$-ADR measure is uniformly rectifiable if it satisfies a version of estimate (1.1) where $\beta_{\mu}$ measures the distance to an optimal vertical hyperplane (Section 1.2). Recent works by Chousionis, Li, and Young provide context for Theorem 1.4 and the importance of the condition (1.1) in the first Heisenberg group. In [12], the authors construct an intrinsic Lipschitz graph (i.e., a graph which is defined with respect to the exterior cone condition in the Heisenberg metric) on which the Riesz transforms associated to fundamental solution of the sub-Laplacian is not $L^{2}(\mu)$ bounded. They also show that the Lipschitz constant of a given intrinsic Lipschitz graph cannot control the constant of the right-hand side of inequality (1.1) with the exponent $2$ replaced by any $p\in[2,4)$. In [14], they show that any given intrinsic Lipschitz graph in $\mathbb{H}$ satisfies (1.1) with exponent $2$ replaced by $4$, with the constant on the right-hand-side depending only the Lipschitz constant of the graph. Both of these results are proved using techniques developed in [46], which we discuss below. This is in contrast to higher Heisenberg groups $\mathbb{H}^{n}$, where, in [13], the same authors show that intrinsic Lipschitz graphs automatically satisfy an estimate like (1.1). Thus, in the presence of the Ahlfors regularity condition, it is natural to seek “$1$-Heisenberg variants” of (1.1) from hypotheses involving boundedness of singular integrals.
The behavior of singular integrals on the Heisenberg group have been heavily studied, with important foundational work by Chousionis and Mattila [15]. Significant results regarding sufficient conditions for the boundedness of all CZOs have been developed in Chousionis, Fässler and Orponen [11, 25]. In [12, 13], Chousionis, Li and Young conjecture that uniform rectifiability is sufficient for the boundedness of all CZOs (Theorem 1.4 provides the necessity) – we do not consider this interesting question here.
In terms of prior work concerning Theorem 1.4, Orponen [51] had previously verified a weaker conclusion that for every $\mathbf{X}\in\operatorname{supp}(\mu)$ and $R>0$,
$\int_{0}^{R}\mathcal{H}^{3}(\{\mathbf{Y}\in\operatorname{supp}(\mu):\beta_{\mu}(\mathbf{Y},r)\geq\varepsilon\})\frac{dr}{r}\leq CR^{3}.$
@@
The Nazarov-Tolsa-Volberg theorem ([47]) has had profound consequences for the study of free boundary problems involving elliptic measures, and, relatedly, connections between continuity of elliptic measures and rectifiability (both qualitative and quantitative). We hope that our results will be an important first step toward establishing analogs of the Nazarov-Tolsa-Volberg theorem in parabolic space and the Heisenberg group. We also think that the HSDC condition should play an important role in the theory of caloric and parabolic measures and parabolic rectifiability.
@@
In 1992 Bishop ([5, Conjecture 10]) conjectured that absolute continuity of harmonic measure with respect to surface measure should imply rectifiability of the boundary. Bishop’s conjecture was solved in 2015 in [1] where the authors effectively used estimates on the Green function to control the Riesz transform and verify the hypotheses of the codimension-one David-Semmes conjecture. Later, in [44], Mourgoglou and Tolsa quantified the main result of [1] using extremely technical arguments involving a corona decomposition with respect to harmonic measure. In [26], Garnett, Mourgoglou, and Tolsa show that, under the codimension-one Ahlfors–regularity condition, if one also assumes either “Carleson measure estimates for harmonic functions” or $\varepsilon$-approximability”—two analytic conditions closely related to the absolute continuity of harmonic measure— then one can conclude uniform rectifiability by establishing that the associated Riesz transforms are $L^{2}$-bounded.
@@
“Two-phase” free boundary problems for harmonic (or, more generally, elliptic measures) are problems where one assumes that two harmonic measures in distinct connected components of a domain enjoy some mutual absolute continuity, in order to conclude that the boundary of the domain enjoys some regularity. In [5], Bishop also conjectured that mutual absolute continuity of harmonic measures in separated connected components of a domain should imply rectifiability of the common boundary of those components. Two-phase problems are more delicate than one phase problems, but similar (and more difficult) arguments can be used to tackle them. In particular, Girela-Sarrión and Tolsa proved a more delicate version of the codimension-one David-Semmes problem in [28] which is suitable for applications to two-phase problems. The main result of [28] was later used in solutions to qualitative and quantitative two-phase problems harmonic measure, we refer the reader to [3], [4], and [2] (the first two authored by Azzam, Mourgoglou, and Tolsa, and the last authored by Azzam, Mourgoglou, Tolsa, and Volberg). In particular, [2] resolves the “two-phase” conjecture of Bishop ([5]). Very recently, Merlo, Mourgoglou, and Puliatti ([41]) have extended many of the results above (one-phase problems, two-phase problems, and the result of Girela-Sarrión and Tolsa) to elliptic measures with coefficients satisfying a “Dini-mean oscillation” condition. We also record that, prior to [41], Puliatti ([53]) had extended the result of Girela-Sarrión and Tolsa to elliptic equations associated to matrices with Hölder coefficients.
@@
In the parabolic case, a converse to Theorem 1.2 is known. In [7, 6, 8] Bortz, Hofmann, Luna-García and Nyström, together with the first author of this paper, developed several foundational quantitative tools for the theory of parabolic quantitative rectifiability. In particular, in [6] and [8] together imply the following
@@
Suppose $\mu$ is a codimension-one parabolic uniformly rectifiable measure on $\mathbb{R}^{n+1}$. Then, if $T$ is a singular integral operator associated to a kernel $K:\mathbb{R}^{n+1}\setminus\{\mathbf{0}\}\rightarrow\mathbb{R}$ such that
@@
We remark that graphs of parabolic Lipschitz functions need not be parabolic uniformly rectifiable (see [7] for a counter-example). This is in contrast to the Euclidean setting, where Lipschitz graphs are a prototypical example of a uniformly rectifiable set. Very recently, the authors have extended their result to the case when $\Gamma$ is a time-symmetric ADR set [^2] Time-symmetric ADR is a stronger condition than the $\bf HSDC$}, see Corollary 1.9. The time symmetric ADR condition is a condition that forces every surface ball to have ample intersection with the set in forward and backward in time.
Theorem 1.7 ([9]).
@@
The proofs of the two theorems above do not rely on analysis of the Riesz transform.
In the past decade there have been some extraordinary results concerning quantitative rectifiability and Heisenberg groups. In [45], Naor and Young resolved several open problems by proving a discrete “vertical versus horizontal” isoperimetric inequality in a discrete Heisenberg group $\mathbb{H}^{n}_{\mathbb{Z}}$, $n\geq 2$. In order to prove this inequality, they prove a corresponding inequality in the continuous Heisenberg group $\mathbb{H}^{n}$, and use the continuous case to deduce the discrete case. The methods of their proof utilize many tools from the area of quantitative rectifiability, in particular the “corona decomposition” type arguments introduced by David and Semmes in [20]. This discrete isoperimetric inequality is used to find an optimal (up to a constant) lower bound on the integrality gap of the Goemans–Linial semidefinite program for the Sparsest Cut Problem, and also establish a lower bound on the distortion of any embedding from the ball of radius $r$ centered at the origin in $(\mathbb{H}_{\mathbb{Z}}^{2},d_{W})$ (the second discrete Heisenberg group equipped with the “word metric”) is at least $C\sqrt{r}$. In [45], Naor and Young prove a “vertical-versus-horizontal” Poincaré inequality in the first Heisenberg group. As a corollary, they obtain, in the first Heisenberg group, bounds on the distortion of embeddings of balls centered at the origin. In this work they develop a new multiscale decomposition for intrinsic Lipschitz graphs, which was later modified in the aforementioned works [12, 13].
There has also been significant advances in extending Preiss’ theorem to the Heisenberg group and parabolic space. These have been developed in the codimension-one setting by Merlo ([42]), and Merlo, Mourgoglou, and Puliatti ([41]), respectively.